News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 9 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours. ACIS door opened nominally on 8/8/99 at approximately 7:00 PM EDT.

Successfully completed IPS deactivation.

The optical cavity pressure was 8.1E-06 Torr. Pressure remains the same during the last 24 hours. Because of the extended time required for pressure decay, no more attempts to turn on HRC high voltage are planned until after first light.

Events scheduled for the next 24 hours
  • Initiate ACIS background measurement calibration.
  • Install on-orbit software loads for OBC A&B.
  • Begin ACA full field search evaluation and FID light characterization.
  • (Gyro Low Rate bias determination was moved downstream 24 hrs to improve overall timeline efficiency.)






  • Return to Status Reports